OpenStudy (anonymous):

Solar radiation of 15000 kcal/m2 yr reaches a meadon. Producers uses 80% of their energy in cellular respiration and supply a net primary productivity of 30000 cal/m2 over a year. what if the efficiency of photosynthesis in this region?

OpenStudy (anonymous):

please help, very confused. there are many examples that are like this on my practice worksheet. and i want to figure out how to do it so i can do the rest

OpenStudy (anonymous):

the answer is 1% but how do you get it?

OpenStudy (koikkara):

Given, Solar radiation= \( 15000 kcal/m^2\) \(80⁄100\) Energy, productivity = \(30000 cal/m^2\)

OpenStudy (koikkara):

@lordofpens formula ?

OpenStudy (anonymous):

useful energy divided by input energy so usually, to find photosynthesis eff i would do GPP/15000 but i dont know the GPP

OpenStudy (koikkara):

. > Gross primary production (GPP) = All CO2 fixed by the plant in photosynthesis > Respiration = R R = CO2 lost from metabolic activity Rp = Respiration by Plants Rh = Respiration by Heterotrophs Rd = Respiration by Decomposers >Net Primary Production = NPP = GPP - Rp >Net Ecosystem Production = NEP = GPP - (Rp + Rh + Rd) @lordofpens

OpenStudy (koikkara):

here, NPP= 30,000 GPP= be X Rp = 80% (Energy) Find, GPP ?? where, NPP = GPP - Rp